OpenStudy (anonymous):

Graph the two equations to solve the system. What is the solution? 1/2 x - 2y = 4 and 4y - x = -8 A. (-2, 0) B. no solution C. (3, 4) D. infinitely many solutions

OpenStudy (anonymous):

well lets see, 1/2x - 2y = 4 => y = 1/4x -2 4y - x = -8 => y = 1/4x -2 if they are the same equation, what is the answer?